Two number cubes have the digits 1, 2, 3, 4, 5, 6 on each. If the number cubes are rolled, what is the probability of the sum being 5 or 8?
Answer by stanbon(75887)

Two number cubes have the digits 1, 2, 3, 4, 5, 6 on each. If the number cubes are rolled, what is the probability of the sum being 5 or 8?
----------------
sum of 5: 1/4 ; 2/3 ; 3/2 ; 4/1
----
sum of 8: 2/6 ; 3/5 ; 4/4 ; 5/3 ; 5/2
========================================
# of ways to succeed: 9
# of possible sums: 36
----
P(5 or 8) = 9/36 = 1/4
==========================
Cheers,
Stan H.
=================